One type of uranium has radioactive decay rate of 0.5%. If 20 pounds of this uranium is available today, find how much will still remain after 30days.

-0.005t
Use y=20(2.7) and let t be 30 days.

Quantity left after 30 days =" "pounds.( round to the nearest tenth)

Answer by stanbon(75887)

One type of uranium has radioactive decay rate of 0.5%. If 20 pounds of this uranium is available today, find how much will still remain after 30days.
----------------------------
A(t) = Ao (1-0.005)^t
A(30) = 20(0.995)^30
A(30) = 20*0.8604
A(30) = 17.208 lbs.
==================
Cheers,
Stan H.